# A Casimir-like probe for 4D Einstein-Gauss-Bonnet gravity

### v1

@@ Abstract
Virtual transitions in a Casimir-like configuration are utilized to probe quantum aspects of the recently proposed four-dimensional Einstein-Gauss-Bonnet (4D EGB) gravity. This study employs a quantum optics-based approach, wherein an Unruh-DeWitt detector (modeled as a two-level atom) follows a radial timelike geodesic, falling freely into an uncharged, nonrotating black hole described by 4D EGB gravity, becoming thermalized in the usual Unruh manner. The black hole, asymptotically Minkowskian, is enclosed by a Casimir boundary proximate to its horizon, serving as a source for accelerated field modes that interact with the infalling detector. Observations are conducted by an asymptotic infinity observer, assuming a Boulware field state. Our numerical analysis reveals that, unlike in Einstein gravity, black holes in 4D EGB gravity can either enhance or suppress the intensity of acceleration radiation, contingent upon the Gauss-Bonnet coupling parameter $\alpha$. Specifically, we observe radiation enhancement for negative $\alpha$ and suppression for positive $\alpha$. These findings offer substantial insights into quantifying the influence of higher-curvature contributions on the behavior of quantum fields in black hole geometries within a 4D spacetime.
@@ I Introduction
The structure of the paper is as follows. The next Sec. II introduces the basics of 4D EGB black hole geometry, accompanied by discussions on the wave equation and the vacuum field state. In Sec. III, we compute the excitation probability or the detector response function of the falling detector. Sec. IV explores possible interpretations of our numerical findings. Finally, conclusions are drawn in Sec. V.
II Conceptual aspects: Our spacetime geometry and the choice of field modes
@@
The static, spherically symmetric metric of an uncharged and nonrotating black hole in 4D EGB gravity is given by Fernandes et al. (2022)
@@
where [^1] We use natural units $c=G=\hbar=1$ throughout.
@@
$r_{\pm}=M^{2}\pm\sqrt{M^{2}-\alpha},$ (4)
of which the one with plus sign is the real exterior horizon of the black hole. Thus, our event horizon is located at $r_{\rm g}=r_{\pm}=M^{2}+\sqrt{M^{2}-\alpha}$. The parameter $\alpha$ can take both positive and negative values within the range $-32M^{2}\leq\alpha\leq 4M^{2}$, as indicated in Refs. Guo and Li (2020); Konoplya and Zinhailo (2020) (also see Glavan and Lin (2020)). It is evident that a positive GB coupling constant $\alpha$ decreases the black hole horizon radius, whereas a negative $\alpha$ increases it. The limit $\alpha=0$ corresponds to the Schwarzschild black hole in GR. These relationships are illustrated graphically in Fig. 1.
@@
II.1 Detector trajectories
@@
Here, $\tau_{0}$ serves as an integration constant, the insignificance of which we establish for the final detector response, as detailed in Sec. III. However, the complexity of the integral for $t$ precludes straightforward analytical computation. Consequently, we resort to numerical methods and present the outcomes in Sec. III. Fig. 2 illustrates the plots of $\tau$ and $t$.
Figure 2: The radial dependence of coordinate time $t$ and proper time $\tau$ illustrates the influence of the Gauss-Bonnet coupling parameter $\alpha$. Panels (a) and (c) depict the proper time for positive and negative $\alpha$, respectively. Similarly, panels (b) and (d) show the coordinate time for positive and negative $\alpha$, respectively.
The plots clearly illustrate that $t$ and $\tau$ exhibit typical Schwarzschild-like behavior. Specifically, $t$, which represents the time measured by an asymptotic observer, diverges as the detector approaches the black hole horizon, located at zero on the rescaled radial coordinate $r-r_{\rm g}$. This divergence signifies that, from the perspective of this observer, the detector never actually crosses the horizon. In contrast, $\tau$ remains finite at the horizon $r-r_{\rm g}$, indicating that from the detector’s own frame of reference, it crosses the horizon in a finite amount of proper time. This disparity highlights the causal structure of black hole horizons and is recognized as gravitational time dilation.
Furthermore, in 4D EGB gravity, the coupling parameter $\alpha$ influences the behavior of $t$ and $\tau$. For positive $\alpha$, which reduces the black hole size as discussed in Sec. II, it takes longer for the detector to approach the horizon as $\alpha$ increases. Conversely, for negative $\alpha$, which inflates the black hole size, the situation is reversed.
II.2 Defining the vacuum state
The response function, or excitation probability, to be calculated in Sec. III, quantifies the detector-field coupling. To achieve this, we must obtain the appropriate field mode by solving the wave equation on the specified spacetime background. Here, we consider the simplest test field: a massless spin-0 Klein-Gordon field, minimally coupled to the spacetime geometry, described by $\nabla_{\mu}\nabla^{\mu}\Phi=0$ Birrell and Davies (1984). Given the spherical symmetry of the spacetime and the presence of a timelike Killing vector $\partial_{t}$, we have $\Phi=\frac{1}{r}Y_{l}(\theta,\phi)\psi(t,r)$, with $Y_{l}$ denoting spherical harmonics and $l$ representing the multipole number. The radial part of the solution, after neglecting the angular dependence $(l=0)$, satisfies the following Schrödinger-like wave equation
@@ II.2 Defining the vacuum state
where we utilized Eq. (2). Additionally, $V(r)$ represents the effective potential experienced by the field, often describing scattering effects in black hole spacetimes Futterman et al. (2012). However, given our focus on the simplest scenario possible, as also demonstrated in Refs. Scully et al. (2018); Bukhari et al. (2023), $V(r)$ can be neglected. One approach to achieve this is by assuming that the frequency $\nu$ of the field mode is sufficiently large, enabling it to surmount the potential barrier imposed by the spacetime. Consequently, the field mode simplifies to
@@ III Detector Response
$\Gamma_{\rm exc}$ $=g^{2}\left|\int\limits_{\infty}^{r_{\rm g}}\mathrm{d}r\,\exp{\left[i\nu\left\{t(r)-r_{*}(r)\right\}\right]}\frac{1}{\sqrt{\frac{r^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r^{3}}}-1\right)}}\exp{\left[i\omega\left\{\frac{2r\sqrt{\sqrt{1+\frac{8\alpha M}{r^{3}}}-1}\tan^{-1}\left(\frac{\sqrt{\sqrt{1+\frac{8\alpha M}{r^{3}}}-1}}{\sqrt{2}}\right)-2\sqrt{2}r}{3\sqrt{\frac{r^{2}\left(\sqrt{1+\frac{8\alpha M}{r^{3}}}-1\right)}{\alpha}}}\right\}\right]}\right|^{2},$ (17)
which results in a complex expression involving nested integrals with respect to $t(r)$ and $r_{*}$. It’s important to note that the limits of integration correspond to the detector’s trajectory from $r=\infty$ to $r=r_{\rm g}$, the horizon of the black hole. Thus, from Eqs. (8) and (12), we derive:
$t(r)=-\int_{\infty}^{r_{\rm g}}\frac{\mathrm{d}r\,}{\left[1+\frac{r^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r^{3}}}\right)\right]\sqrt{\frac{r^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r^{3}}}-1\right)}},\ \ \ \ r_{*}$ $=\int_{r_{\rm g}}^{\infty}\frac{\mathrm{d}r\,}{1+\frac{r^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r^{3}}}\right)}.$ (18)
Consider now the substitution $r=r_{\rm g}z$, where $\mathrm{d}r=r_{\rm g}\mathrm{d}z$. Using this transformation of variables, we may rewrite $t(r)$ in Eq. (18) as follows:
$t(z)=-\int_{\infty}^{1}\frac{\mathrm{d}z\,r_{\rm g}}{\left[1+\frac{r_{\rm g}^{2}z^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r_{\rm g}^{3}z^{3}}}\right)\right]\sqrt{\frac{r_{\rm g}^{2}z^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r_{\rm g}^{3}z^{3}}}-1\right)}}.$ (19)
@@
$t(x)=\int_{0}^{\infty}\frac{\mathrm{d}x\,r_{\rm g}}{\left[1+\frac{r_{\rm g}^{2}(x+1)^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r_{\rm g}^{3}(x+1)^{3}}}\right)\right]\sqrt{\frac{r_{\rm g}^{2}(x+1)^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r_{\rm g}^{3}(x+1)^{3}}}-1\right)}}.$ (20)
@@
$r_{*}(x)$ $=\int_{0}^{\infty}\frac{\mathrm{d}x\,r_{\rm g}}{1+\frac{\left[r_{\rm g}(x+1)\right]^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{\left[r_{\rm g}(x+1)\right]^{3}}}\right)}.$ (21)
@@
$\Gamma_{\rm exc}$ $=g^{2}r_{\rm g}^{2}\left|\int_{0}^{\infty}\mathrm{d}x\ \ \ \exp{\left[i\nu\left\{t(x)-r_{*}(x)\right\}\right]}\frac{\mathcal{G}}{\sqrt{\frac{(r_{\rm g}[x+1])^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{(r_{\rm g}\left[x+1\right])^{3}}}-1\right)}}\right|^{2},$
@@
$\mathcal{G}=\exp{\left[i\omega\left\{\frac{2r_{\rm g}\left(x+1\right)\sqrt{\sqrt{1+\frac{8\alpha M}{(r_{\rm g}\left[x+1\right])^{3}}}-1}\tan^{-1}\left(\frac{\sqrt{\sqrt{1+\frac{8\alpha M}{(r_{\rm g}\left[x+1\right])^{3}}}-1}}{\sqrt{2}}\right)-2\sqrt{2}\left[r_{\rm g}(x+1)\right]}{3\sqrt{\frac{(r_{\rm g}\left[x+1\right])^{2}\left(\sqrt{1+\frac{8\alpha M}{(r_{\rm g}\left[x+1\right])^{3}}}-1\right)}{\alpha}}}\right\}\right]}.$ (22)
Figure 3: Impact of Gauss-Bonnet coupling $\alpha$ on detector response $\Gamma_{exc}$:(a) positive $\alpha$, and (b) negative $\alpha$, for a fixed black hole mass $M$.(c) indicates impact of detector transition frequency $\omega$, and (d) is the log-log plot of $\Gamma_{\rm exc}$ to emphasize the BE-type nature of the spectrum. We chose $M=1$, $\omega=50$ and $g=1$.

### v2

@@ Abstract
Virtual transitions in a Casimir-like configuration are utilized to probe quantum aspects of four-dimensional Einstein-Gauss-Bonnet (4D EGB) gravity. This study employs a quantum optics-based approach, wherein an Unruh-DeWitt detector (modeled as a two-level atom) follows a radial timelike geodesic, falling freely into an uncharged, nonrotating black hole described by 4D EGB gravity, becoming thermalized in the usual Unruh manner. The black hole, asymptotically Minkowskian, is enclosed by a Casimir boundary proximate to its horizon, serving as a source for accelerated field modes that interact with the infalling detector. Observations are conducted by an asymptotic infinity observer, assuming a Boulware field state. Our numerical analysis reveals that, unlike in Einstein gravity, black holes in 4D EGB gravity can either enhance or suppress the intensity of acceleration radiation, contingent upon the Gauss-Bonnet coupling parameter $\alpha$. Specifically, we observe radiation enhancement for negative $\alpha$ and suppression for positive $\alpha$. These findings offer substantial insights into quantifying the influence of higher-curvature contributions on the behavior of quantum fields in black hole geometries within a 4D spacetime.
@@ I Introduction
The structure of the paper is as follows. The next Sec. III introduces the basics of 4D EGB black hole geometry, accompanied by discussions on the wave equation and the vacuum field state. In Sec. IV, we compute the excitation probability or the detector response function of the falling detector. Sec. V explores possible interpretations of our numerical findings. Finally, conclusions are drawn in Sec. VII.
II Horizon Brightened acceleration radiation: An outlook
The core idea behind radiation emission from the vacuum in the typical Hawking-Unruh paradigm is the fact that vacuum fluctuations do not follow Poincaré invariance in presence of external influences, such as gravity or acceleration Birrell and Davies (1984). Hence it is tantamount to say that quantum vacuum gets polarized due to these factors. Beyond famous Hawking-Unruh radiation, this produces myriad of non-stationary QED effects Masood A. S. Bukhari and Wang (2024); Nation et al. (2012), including dynamical Casimir effect as mentioned in Introduction. Out of many interesting mechanisms, one possible idea is to utilize moving boundaries. In principle, a moving boundary (or mirror) potentially alters the structure of vacuum, resulting in the creation and annihilation of radiation quanta Haro and Elizalde (2006). These models are very crucial for understanding particle production in many cosmological scenarios and radiation emitted from black holesMasood A. S. Bukhari and Wang (2024). Our investigation in the present work revolving around Ref.Scully et al. (2018) is grounded in the concept of these moving mirrors or accelerated boundaries.
Figure 1: Schematic view of acceleration radiation emitted from a two-level atom falling into a black hole. The Casimir boundary provides accelerated field modes with which the atom interacts with, besides shielding it from Hawking radiation. Hence, no quanta in the accelerated radiation has any Hawking contribution.
II.1 The Model
To best comprehend the workings of our model and subsequent investigations, we first provide a pictorial representation in Fig. 1 to put things as easy as possible. With that, we believe it captures the essence of the setup introduced in Scully et al. (2018).
The configuration assumes a black hole with its event horizon as indicated in Fig.1 above which hovers one end (mirror or Casimir boundary) of a cavity [^1] Consider the mirror being held in position firmly against black hole’s gravity by some rockets! whose other end is assumed to be fixed somewhere at asymptotic infinity. A two-level atom acting as a typical Unruh-De Witt detector falls freely from infinity toward the event horizon of the black hole. Thus, in accordance with the equivalence principle of general relativity, atom is in an inertial frame of reference (zero four-acceleration), while the Casimir boundary is in an accelerated frame of reference. An observer, located at infinity, defines the field around the black hole in its vacuum state. Specifically, as such observer is at asymptotically flat region, the appropriate vacuum for such an observer is the Boulware state Birrell and Davies (1984). Since atom is in inertial frame, it is not expected to emit any radiation, and this the essence of quantum field theory in curved spaces Birrell and Davies (1984). However, as argued in Scully et al. (2018), the vacuum field modes interacting with the mirror are polarized in a similar manner to that of Hawking radiation. Thus there is a relative acceleration between atom and the Casimir boundary which gives rise to the acceleration radiation detected by the asymptotic observer. The Casimir boundary, in addition to providing accelerating field modes, also shields the atom from any Hawking quanta (shown as orange colored ring). The boundary can be envisioned by either assuming a reflecting mirror as shown in Fig.1 or an enclosing spherical surface shrouding the entire black hole Bukhari et al. (2023); Bukhari and Wang (2024).
II.2 The Significance
It is quite intriguing to see from above that normally, such observers with Boulware vacuum state do not observe any Hawking quanta. This radiation emission thus has purely a quantum optical origin and has surprising similarities to that of Hawking radiation. The radiation has a thermal Planck distribution with a characteristic temperature and gives rise to an entropy-area correspondence. While defining the associated entropy, it becomes crucial to consider area lost by the black hole due to the radiation. From the relation Scully et al. (2018),
$\dot{S}_{\rm HBAR}=\frac{k_{\rm B}c^{3}}{4\hbar G}\dot{A},$ (1)
where $k_{\rm B},c,\hbar$ and $G$ represent Boltzmann constant, velocity of light, Planck’s constant Newtons’ gravitational constant, respectively, one can see the rate of change of the entropy being related to the rate of loss of area $A$ of the black hole through HBAR flux. Despite these similarities between Hawking radiation and HBAR, certain parameters like phase correlations between quanta make HBAR different from Hawking emission, avoiding any information loss paradox.
Above setup undoubtedly provides new insights into Hawking-Unruh effect and equivalence principle by assuming some kind of atom-mirror symmetry Fulling and Wilson (2019); Svidzinsky et al. (2018). However, certain extreme scenarios in quantum gravity, where the field satisfies modified dispersion relation Kempf et al. (1995), might indicate limited regimes of applicability of HBAR emission Chatterjee et al. (2021).
III Our spacetime geometry and the choice of field modes
In Einstein gravity, the 4D Einstein-Hilbert action reads
$S_{\rm EH}=\int\mathrm{d}^{4}x\sqrt{-g}R,$ (2)
where $R$ denotes Ricci curvature, and $g$ is the determinant of metric tensor $g_{\mu\nu}$. According to Lovelock theorem Lovelock (1971, 1972); Lanczos (1938), GR is the unique gravitational theory in four dimensions because it satisfies the criteria of being diffeomorphism-invariant, possesses metricity, and its equations of motion are all second-order. In higher spacetime dimensions, GB action
$S_{\rm GB}=\int\mathrm{d}^{D}x\sqrt{-g}\left(R+\alpha\mathcal{G}\right)$ (3)
satisfies these conditions also. Here, $\mathcal{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ is the GB invariant term, $R^{\mu\nu}$ and $R^{\mu\nu\rho\sigma}$ denote the Ricci and Riemann tensors, respectively. If one varies Eq. (3) with respect to $g_{\mu\nu}$, one gets gravitational field equations
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}+\alpha H_{\mu\nu}=T_{\mu\nu},$ (4)
where we wrote
$H_{\mu\nu}=$ $2RR^{\mu\nu}-4R_{\mu\sigma}R^{\sigma}_{\nu}-4R_{\mu\sigma\nu\rho}R^{\sigma\rho}$ (5)
$-2R_{\mu\sigma\nu\delta}R^{\sigma\rho\delta}_{\nu}-\frac{1}{2}\mathcal{G}g_{\mu\nu},$ (6)
and $T_{\mu\nu}$ represents the energy-momentum tensor.
@@ III Our spacetime geometry and the choice of field modes
The static, spherically symmetric metric of an uncharged and nonrotating black hole in 4D EGB gravity gotten from Eq. 4 is given by Fernandes et al. (2022)
@@
where [^2] We use natural units $c=G=\hbar=k_{\rm B}=1$ throughout from now on.
@@
$r_{\pm}=M\pm\sqrt{M^{2}-\alpha},$ (10)
of which the one with plus sign is the (outer) horizon of the black hole. The parameter $\alpha$ can take both positive and negative values, as indicated in Refs. Guo and Li (2020); Konoplya and Zinhailo (2020) (also see Glavan and Lin (2020)). Moreover, for $\alpha>0$, there is a minimum mass of the black hole given by
$M_{\rm min}=\sqrt{\alpha},$ (11)
below which $f(r)$ does not vanish and one encounters naked singularities. The negative values of $\alpha$ have been shown to be tightly constrained Charmousis et al. (2022); Clifton et al. (2020). In this paper, working in natural units, we take range of $\alpha$ as $[-1,1]$, consistent with Eq.10. If one likes to compute parameters in dimensional units, then one notes that mass $M$ in Eq. 9 has to be replaced by $GM/c^{2}$, which in view of Eq. 10 necessitates $\alpha<G^{2}M^{2}/c^{4}$ to avoid any naked singularity. The bounds for values of $\alpha$ as noted in Charmousis et al. (2022) for the case of black holes vis-à-vis gravitational wave event GW190814190814190814190814 Abbott et al. (2020) are $-10^{-30}\rm m^{2}\leq\alpha<5.9\times 10^{7}\rm m^{2}$, which are tighter (especially for negative $\alpha$) than the bounds provided in Ref.Clifton et al. (2020) with $-10^{10}<\alpha<10^{9}\rm m^{2}$. The range $\alpha\in[-1,1]$ we consider in this work is in natural units which makes them arbitrary and hence suitable for any constraints and system of units. However, all calculations done in this paper should be viewed in compliance with the bounds provided in Refs. Charmousis et al. (2022); Clifton et al. (2020).
It is evident that a positive GB coupling constant $\alpha$ decreases the black hole horizon radius, whereas a negative $\alpha$ increases it. The limit $\alpha=0$ corresponds to the Schwarzschild black hole in GR. These relationships are illustrated graphically in Fig. 2.
@@
Although $4$D EGB gravity has many interesting phenomenological predictions compared to GR, its overall structure and consistency has been point of major debate among researchers. For example, the tensor $H_{\rm\mu\nu}$ given by Eq.(5) has been shown not to possess an inherently $4$D description in terms of a covariantly-conserved rank-two tensor in $4$D Gürses et al. (2020); Gurses et al. (2020); Arrechea et al. (2021, 2020). Moreover, the possibility of a nontrivial D4absent4$\rightarrow$ 4→ 4 limit of Eq.3 has also been investigated using tree-level graviton scattering amplitudes Bonifacio et al. (2020). It has been found that $4$D EGB gravity in some sense can be understood as a description arising out of a scalar field, hence making the theory a part of scalar-tensor theories Bonifacio et al. (2020); Fernandes et al. (2020). Furthermore, the divergences in the on-shell action Mahapatra (2020) makes it impossible to variationallly complete field equations Hohmann et al. (2021). In view of these concerns and criticisms, the regularization of $4$D EGB has been a subject of intense investigation. For a detailed outlook, we refer the interested reader to the Refs.Fernandes et al. (2020); Fernandes (2021); Hennigar et al. (2020); Kobayashi (2020); Lu and Pang (2020); Aoki et al. (2020).
III.1 Detector trajectories
@@
Here, $\tau_{0}$ serves as an integration constant, the insignificance of which we establish for the final detector response, as detailed in Sec. IV. However, the complexity of the integral for $t$ precludes straightforward analytical computation. Consequently, we resort to numerical methods and present the outcomes in Sec. IV. Fig. 3 illustrates the plots of $\tau$ and $t$.
Figure 3: The radial dependence of coordinate time $t$ and proper time $\tau$ illustrates the influence of the Gauss-Bonnet coupling parameter $\alpha$. Panels (a) and (c) depict the proper time for positive and negative $\alpha$, respectively. Similarly, panels (b) and (d) show the coordinate time for positive and negative $\alpha$, respectively. For all plots, we set black hole mass $M=1$.
The plots clearly illustrate that $t$ and $\tau$ exhibit typical Schwarzschild-like behavior. Specifically, $t$, which represents the time measured by an asymptotic observer, diverges as the detector approaches the black hole horizon, located at zero on the rescaled radial coordinate $r-r_{+}$. This divergence signifies that, from the perspective of this observer, the detector never actually crosses the horizon. In contrast, $\tau$ remains finite at the horizon $r-r_{+}$, indicating that from the detector’s own frame of reference, it crosses the horizon in a finite amount of proper time. This disparity highlights the causal structure of black hole horizons and is recognized as gravitational time dilation.
Furthermore, in 4D EGB gravity, the coupling parameter $\alpha$ influences the behavior of $t$ and $\tau$. For positive $\alpha$, which reduces the black hole size as discussed in Sec. III, it takes longer for the detector to approach the horizon as $\alpha$ increases. Conversely, for negative $\alpha$, which inflates the black hole size, the situation is reversed.
III.2 Defining the vacuum state
The response function, or excitation probability, to be calculated in Sec. IV, quantifies the detector-field coupling. To achieve this, we must obtain the appropriate field mode by solving the wave equation on the specified spacetime background. Here, we consider the simplest test field: a massless spin-0 Klein-Gordon field, minimally coupled to the spacetime geometry, described by $\nabla_{\mu}\nabla^{\mu}\Phi=0$ Birrell and Davies (1984). Given the spherical symmetry of the spacetime and the presence of a timelike Killing vector $\partial_{t}$, we have $\Phi=\frac{1}{r}Y_{l}(\theta,\phi)\psi(t,r)$, with $Y_{l}$ denoting spherical harmonics and $l$ representing the multipole number. The radial part of the solution, after neglecting the angular dependence $(l=0)$, satisfies the following Schrödinger-like wave equation
@@ III.2 Defining the vacuum state
where we utilized Eq. (8). Additionally, $V(r)$ represents the effective potential experienced by the field, often describing scattering effects in black hole spacetimes Futterman et al. (2012). However, given our focus on the simplest scenario possible, as also demonstrated in Refs. Scully et al. (2018); Bukhari et al. (2023), $V(r)$ can be neglected. One approach to achieve this is by assuming that the frequency $\nu$ of the field mode is sufficiently large, enabling it to overcome the potential barrier imposed by the spacetime. Consequently, the field mode simplifies to
@@ IV Detector Response
$\Gamma_{\rm exc}$ $=g^{2}\left|\int\limits_{\infty}^{r_{+}}\mathrm{d}r\,\exp{\left[i\nu\left\{t(r)-r_{*}(r)\right\}\right]}\frac{1}{\sqrt{\frac{r^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r^{3}}}-1\right)}}\exp{\left[i\omega\left\{\frac{2r\sqrt{\sqrt{1+\frac{8\alpha M}{r^{3}}}-1}\tan^{-1}\left(\frac{\sqrt{\sqrt{1+\frac{8\alpha M}{r^{3}}}-1}}{\sqrt{2}}\right)-2\sqrt{2}r}{3\sqrt{\frac{r^{2}\left(\sqrt{1+\frac{8\alpha M}{r^{3}}}-1\right)}{\alpha}}}\right\}\right]}\right|^{2},$ (24)
which results in a complex expression involving nested integrals with respect to $t(r)$ and $r_{*}$. It’s important to note that the limits of integration correspond to the detector’s trajectory from $r=\infty$ to $r=r_{+}$, the horizon of the black hole. Thus, from Eqs. (15) and (19), we derive:
$t(r)=-\int_{\infty}^{r_{+}}\frac{\mathrm{d}r\,}{\left[1+\frac{r^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r^{3}}}\right)\right]\sqrt{\frac{r^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r^{3}}}-1\right)}},\ \ \ \ r_{*}$ $=\int_{r_{+}}^{\infty}\frac{\mathrm{d}r\,}{1+\frac{r^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r^{3}}}\right)}.$ (25)
Consider now the substitution $r=r_{+}z$, where $\mathrm{d}r=r_{+}\mathrm{d}z$. Using this transformation of variables, we may rewrite $t(r)$ in Eq. (25) as follows:
$t(z)=-\int_{\infty}^{1}\frac{\mathrm{d}z\,r_{+}}{\left[1+\frac{r_{+}^{2}z^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r_{+}^{3}z^{3}}}\right)\right]\sqrt{\frac{r_{+}^{2}z^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r_{+}^{3}z^{3}}}-1\right)}}.$ (26)
@@
$t(x)=\int_{0}^{\infty}\frac{\mathrm{d}x\,r_{+}}{\left[1+\frac{r_{+}^{2}(x+1)^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{r_{+}^{3}(x+1)^{3}}}\right)\right]\sqrt{\frac{r_{+}^{2}(x+1)^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{r_{+}^{3}(x+1)^{3}}}-1\right)}}.$ (27)
@@
$r_{*}(x)$ $=\int_{0}^{\infty}\frac{\mathrm{d}x\,r_{+}}{1+\frac{\left[r_{+}(x+1)\right]^{2}}{2\alpha}\left(1-\sqrt{1+\frac{8\alpha M}{\left[r_{+}(x+1)\right]^{3}}}\right)}.$ (28)
@@
$\Gamma_{\rm exc}$ $=g^{2}r_{+}^{2}\left|\int_{0}^{\infty}\mathrm{d}x\ \ \ \exp{\left[i\nu\left\{t(x)-r_{*}(x)\right\}\right]}\frac{\mathcal{G}}{\sqrt{\frac{(r_{+}[x+1])^{2}}{2\alpha}\left(\sqrt{1+\frac{8\alpha M}{(r_{+}\left[x+1\right])^{3}}}-1\right)}}\right|^{2},$
@@
$\mathcal{G}=\exp{\left[i\omega\left\{\frac{2r_{+}\left(x+1\right)\sqrt{\sqrt{1+\frac{8\alpha M}{(r_{+}\left[x+1\right])^{3}}}-1}\tan^{-1}\left(\frac{\sqrt{\sqrt{1+\frac{8\alpha M}{(r_{+}\left[x+1\right])^{3}}}-1}}{\sqrt{2}}\right)-2\sqrt{2}\left[r_{+}(x+1)\right]}{3\sqrt{\frac{(r_{+}\left[x+1\right])^{2}\left(\sqrt{1+\frac{8\alpha M}{(r_{+}\left[x+1\right])^{3}}}-1\right)}{\alpha}}}\right\}\right]}.$ (29)
Figure 4: Impact of Gauss-Bonnet coupling $\alpha$ on detector response $\Gamma_{\rm exc}$:(a) positive $\alpha$, and (b) negative $\alpha$, for a fixed black hole mass $M$.(c) indicates impact of detector transition frequency $\omega$ where we set $\alpha=0.1$, and (d) is the log-log plot of $\Gamma_{\rm exc}$ to emphasize the BE-type nature of the spectrum. We chose $g=1$ throughout.